A rocket moving upwards with an acceleration of 4g records the weight of a man as 300kgf.what is it's actual weight
 

Grade:10

2 Answers

Arun
4 years ago
Apparant weight of the man = 300 kgwt Acceleration of the rocket =4g
Then, we have,
Net upward force=Net downward force, then
Apparent weight=m(g+a)
300kgwt=m(g+4g)
5mg=300kgwt
mg=60kgwt i.e.,
Actual weight of the man= 60kgwt
Vikas TU
4 years ago
Dear student 
In Generalised term 
Let, ‘m’ be  the mass of the man, ‘g’ be the acceleration due to gravity, ‘a’ is the acceleration of the lift upwards.
Forces on the floor by the man are:
Gravitational force downward = mg
The reaction on the floor due to the upward acceleration of the lift = ma
So, the net force on the floor by the man = mg + ma = m(g + a)
This force m(g + a) is the apparent weight of the man.
Good Luck